# Phase-Field simulations of dynamic wetting by the Volume-Of-Fluid method

### v1

@@ Abstract
In this work, we present a toy model for dynamic contact lines using the Volume-of-Fluid method, designed to match Phase-Field results at a lower computational cost. We use a theoretical relation to model the mass diffusion at the contact line through a simple relation for the contact angle. In practice, the computational boundary is lifted a specific distance away from the wall where we apply a Navier boundary condition coupled with a dynamic apparent contact angle that accounts for the curvature of the interface. The comparison of the model results with respect to the full Phase-Field ones, on two different setups with the same input parameters, demonstrates the robustness of the model. This offers a practical and efficient alternative to solve contact line problems at lower cost.
keywords:
Contact line, Wetting, Phase-Field, Cahn-Hilliard, Volume-Of-Fluid, Navier slip
1 Introduction
@@
The contact line that forms at the intersection of a fluid-fluid and a solid boundary is a key element of many natural and several technological processes. The dynamics of the contact line has been a subject of debate for several decades, see Blake & Shikhmurzaev (2002); Hocking (1992); Eggers & Evans (2004); Wilson et al. (2006); Shikhmurzaev (2006); Bonn (2009); Snoeijer & Andreotti (2013). The classical no-slip boundary condition at a moving fluid interface leads to a force singularity at the contact line, as demonstrated by Huh & Scriven (1971), implying that contact line motion would be impossible.
@@
In the context of continuum mechanics, in sharp-interface models, one common approach is to introduce a small slip velocity at the contact line using the Navier boundary condition. The slip length, typically on the order of nanometers, is often associated with surface roughness or the mean free path of molecules, see Lācis et al. (2020). The Navier slip model combined with an equilibrium contact angle on a mesh-aligned boundary has been extensively studied across various physical setups using different numerical methods. For instance, a rapidly advancing contact line in a curtain coating scenario was simulated in a Volume-of-Fluid (VOF) framework by Fullana et al. (2020), different numerical methods with the same contact line model were tested in a capillary rise setup by Gründing et al. (2020), a withdrawing plate was simulated using the VOF method with a mesh-dependent slip length by Afkhami & Bussmann (2009a). Spreading droplets were explored with mesh-dependent slip by Legendre & Maglio (2015), and the Level-Set method was employed for a spreading drop case, incorporating contact line hysteresis by varying the contact angle based on the contact line speed by Spelt (2005). Recently, a Generalized Navier Boundary Condition implemented within the same sharp interface framework was shown to regularize the logarithmic singularity at the contact line by allowing a deviation of dynamic contact angle with respect to the equilibrium one through the uncompensated Young’s stress, see Kulkarni et al. (2023); Fullana et al. (2024). These models have in common that they require empirical information on the slip length, which can be viewed as a fitting parameter.
@@
On the other hand, in the Phase-Field method (PF) the Cahn-Hilliard equations are formulated from the thermodynamics of an immiscible two-component mixture and a phase function is used to represent the moving interface. The velocity field can satisfy a no-slip condition at the contact line, and the contact line moves due the diffusive mass transport that is present in a two-component system, see Jacqmin (1999); Yue et al. (2010). While the phase-field method has the advantage of needing less empirical fitting, it is considerably more expensive. Although it has been shown that it is possible to match the results obtained on dynamic contact lines by fine-tuning the different parameters of the models, see Lācis et al. (2020, 2022), an ideal model would mimic the mass diffusion by a simple model; that is to obtain results essentially equivalent to the PF ones by doing much cheaper VOF simulations.
@@
2 Sharp interface model
The Volume-Of-Fluid method for representing fluid interfaces coupled with a flow solver is well-known to be suited for solving interfacial flows, see Scardovelli & Zaleski (1999); Popinet & Zaleski (1999); Tryggvason et al. (2011). In the present work, we use the free software Basilisk, a platform for the solution of partial differential equations on adaptive Cartesian meshes, described in Popinet (2003, 2009, 2015). We consider two incompressible, Newtonian and non-miscible fluids. We assume that there is no mass and heat transfer at the fluid-fluid interface, that the surface tension between the two fluids $\sigma$ is constant, and neglect gravity. The two fluids will be denoted l and g for liquid and gas phases, with densities and dynamical viscosities $\rho_{l}$ and $\rho_{g}$, $\mu_{l}$ and $\mu_{g}$ respectively. In the one-fluid formulation, the unsteady, incompressible, variable density, Navier-Stokes equations read
$\dfrac{\partial\rho}{\partial t}+\mathbf{\nabla}\cdot(\rho\,\mathbf{u})$ $=0,$ (2.1)
$\rho\left(\dfrac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\mathbf{\nabla}\mathbf{u}\right)$ $=-\mathbf{\nabla}P+\mathbf{\nabla}\cdot(2\mu\mathbf{D})+\mathbf{F}_{\sigma},$ (2.2)
$\mathbf{\nabla}\cdot\mathbf{u}$ $=0,$ (2.3)
with $\mathbf{u}\equiv\mathbf{u}(\mathbf{x},t)$ the flow velocity, $P\equiv P(\mathbf{x},t)$ the pressure, $\mathbf{D}=(\mathbf{\nabla}\mathbf{u}+\mathbf{\nabla}\mathbf{u}^{T})/2$ the rate of deformation tensor, $\rho\equiv\rho(\mathbf{x},t)$ the density, $\mu\equiv\mu(\mathbf{x},t)$ the viscosity and $\mathbf{F}_{\sigma}$ the surface tension force. The surface tension term is modeled using the continuum surface force of Brackbill et al. (1992) so that $\mathbf{F}_{\sigma}=\sigma\,\kappa\,\nabla\chi$ with $\kappa$ the curvature, $\mathbf{n}$ the normal unit vector to the interface and $\chi$ the Heaviside function
$\chi(\mathbf{x},t)=\left\{\begin{array}[]{rl}0,&\text{in the gas phase},\\ 1,&\text{in the liquid phase},\end{array}\right.$ (2.4)
also used to locate the interface and define the viscosity and density, see Tryggvason et al. (2011). The volume fraction $c(\mathbf{x},t)$ corresponding to the average value of $\chi$ in each cell is used to replace the advection of the density (2.1) by the following advection equation:
$\dfrac{\partial c}{\partial t}+\mathbf{u}\cdot\mathbf{\nabla}c=0.$ (2.5)
@@ 2 Sharp interface model
The incompressible Navier-Stokes equations are approximated using a time-staggered approximate projection method on a Cartesian grid. The advection term is discretized with the explicit and conservative Bell–Colella–Glaz second-order unsplit upwind scheme, see Bell et al. (1989). For the discretization of the viscous diffusion term, a second-order Crank–Nicholson fully-implicit scheme is used. Spatial discretization is achieved using a quadtree (octree in three dimensions) adaptive mesh refinement on collocated grids. The adaptive mesh technique is based on the wavelet decomposition of the variables such as the velocity field or the volume fraction and is efficient and critical to the success of the simulations as it allows high resolution only where needed, therefore reducing the cost of calculation.
@@
The resolution of the surface tension term is directly dependent on the accuracy of the curvature calculation. The Height-Function methodology is a VOF-based technique for calculating interface normals and curvatures, see Afkhami & Bussmann (2008, 2009b). About each interface cell, fluid ‘heights’ are calculated by summing fluid volume in the grid direction closest to the normal of the interface. In two dimensions, a $7\times 3$ stencil around an interface cell is constructed and the heights are evaluated by summing volume fractions horizontally
$h_{j}=\sum_{k=i-3}^{k=i+3}c_{j,k}\Delta_{\text{V}},$ (2.7)
@@
$\begin{array}[]{c}{\mathbf{n}=\left(h_{x},-1\right)},\\ \\ {\kappa=\dfrac{h_{xx}}{\left(1+h_{x}^{2}\right)^{3/2}}},\end{array}$ (2.8)
where $h_{x}$ and $h_{xx}$ are discretized using second-order central differences. This method is also used to model the contact angle at the triple point. The orientation of the interface, characterized by the contact angle – the angle between the normal to the interface at the contact line and the normal to the solid boundary – is imposed in the contact line cell. It is important to note that a numerical specification of the contact angle affects the overall flow calculation in two ways: (i) it defines the orientation of the VOF reconstruction in cells that contain the contact line; (ii) it influences the calculation of the surface tension term by affecting the curvature computed in cells at and near the contact line. The boundary conditions in the VOF toy model, derived in Section 4, will result in a simple slip model coupled to a dynamic contact angle. The slip model, or Navier boundary condition, has been implemented and used in various studies within this numerical framework in Fullana et al. (2020); Lācis et al. (2020). In two dimensions, the Navier boundary condition reads
$\begin{split}u_{x}|_{y=0}+\lambda\dfrac{\partial u_{x}}{\partial y}|_{y=0}&=U,\\ u_{y}|_{y=0}&=0,\end{split}$ (2.9)
with $u_{x}|_{y=0}$ and $u_{y}|_{y=0}$ the $x$ and $y$ component of the velocity at the solid boundary, $\lambda$ the slip length, and $U$ the prescribed velocity of the moving solid. Note that (2.9) is written with the boundary located at $y=0$; the location of the computational boundary will be lifted in the toy model (Section 4) such that the boundary is located at $y=+\lambda$. Moreover, it is important to remark that the accuracy of the numerical method is dependent on the resolution of the smallest length scale, the slip length $\lambda$. In the following numerical setups, the grid size will be chosen such that $\lambda/\Delta_{\text{V}}\simeq 4$.
3 Diffuse interface model
In the Phase-Field method, the governing equations are derived from the thermodynamic potentials of the system, together with the assumption of a surface energy associated with an interface, see Cahn & Hilliard (1958a); Jacqmin (2000); Yue et al. (2010). It is, therefore, possible to consider different physical situations with relative ease. It is also straightforward to implement numerically, since interfaces are not tracked explicitly. Instead, a variable is introduced that has different constant values in the two phases with a steep transition between the two in the diffuse interface, see Jacqmin (2000); Carlson (2012). One of the major drawbacks of this method, however, is that the width of the interface $\epsilon$ must be small to match the proper interface dynamics.
@@ 3 Diffuse interface model
$\dfrac{\partial C}{\partial t}=\nabla\cdot(-\mathbf{F}_{d}-\mathbf{F}_{c}),$ (3.1)
@@
$\mathbf{F}_{c}=\mathbf{u}\cdot\nabla C,$ (3.2)
@@
$\phi=\beta\Psi^{\prime}\left(C\right)-\alpha\nabla^{2}C.$ (3.4)
@@
$\sigma=\dfrac{2}{3}\sqrt{2\alpha\beta},\quad\quad\epsilon=\sqrt{\dfrac{\alpha}{\beta}}.$ (3.5)
@@
$\Psi\left(C\right)=\dfrac{\left(C+1\right)^{2}\left(C-1\right)^{2}}{4}.$ (3.6)
With the definitions above (3.1)- (3.6), the convection-diffusion equation, referred to as the Cahn-Hilliard (1958b) equation, can be written as
$\dfrac{\partial C}{\partial t}=\nabla\cdot\left[M\nabla\left(\beta\Psi^{\prime}\left(C\right)-\alpha\nabla^{2}C\right)\right]-\mathbf{u}\cdot\nabla C.$ (3.7)
Equation (3.7) is the Phase-Field analog of (2.5). The boundary conditions for this fourth-order partial differential equation will be detailed in our study of dynamic wetting by the Phase-Field method. Taking the same notation as in (2.2), the momentum equation in the Navier-Stokes equations becomes
$\rho\left(\partial_{t}\mathbf{u}+\mathbf{u}\cdot\nabla\mathbf{u}\right)=-\nabla p+\nabla\cdot(2\mu\mathbf{D})+\phi\nabla C\,$ (3.8)
@@
$\rho({C})$ $\equiv\rho_{l}\dfrac{C+1}{2}+\rho_{g}\dfrac{C-1}{2},$ (3.9)
@@
We now state the boundary conditions in the presence of a contact line on a solid, characterized by its equilibrium angle (the angle between the fluid-fluid interface and the wall). The convection-diffusion equation (3.7) is a fourth-order partial differential equation and requires two boundary conditions. First, we impose a non-equilibrium wetting condition, see Jacqmin (2000); Qian et al. (2003) on the solid wall,
$-\mu_{f}\epsilon\left(\dfrac{\partial C}{\partial t}+\mathbf{u}\cdot\nabla C\right)=\alpha\nabla C\cdot\mathbf{n}-\sigma\cos\theta_{\mathrm{e}}\>g^{\prime}\left(C\right),$ (3.10)
@@
$g\left(C\right)=\dfrac{1}{4}(2-3C+C^{3}),$ (3.11)
@@
$\nabla\phi\cdot\mathbf{n}=0.$ (3.12)
The Phase-Field equations are written into the weak form, and solved using an open-source finite-element software FreeFEM by Hecht, Pironneau and others (see Hecht, 2012; Pironneau et al., 2020), allowing the specification of a finite-element weak form as input. Linear elements were used for the PF variables, while the fluid flow was resolved using Taylor–Hood elements (quadratic for velocity and linear for pressure). The numerical code is available in the Github repository FreeFEM-NS-CH. Similarly to the VOF implementation, the accuracy of the method depends on the resolution of the smallest length scale of the system, the interface width $\epsilon$. In the PF setups, the grid size will be chosen such that $\epsilon/\Delta_{\text{P}}\simeq 16$.
4 The toy model: relating mass flux to curvature

### v2

# Mass diffusion and bending in dynamic wetting by phase-field and sharp interface models

@@ Abstract
Dynamic wetting poses a well-known challenge in classical sharp interface formulation as the no-slip wall condition leads to a contact line singularity that is typically regularized with a Navier boundary condition, often requiring empirical fitting for the slip length. On the other hand, this paradox does not appear in phase field models as the contact line moves through diffusive mass transport. In this work, we present a toy model that accounts for mass diffusion at the contact line within a sharp interface framework. This model is based on a theoretical relation derived from the Cahn-Hilliard equations, which links the total diffusive mass transport to the curvature at the wall. From an estimate of the chemical potential on a curved interface, we obtain an expression for the width of the highly curved region $\delta$ and the apparent angle. In the sharp interface model, we then introduce a fictitious boundary, displaced of a distance $\delta$ into the domain, where a Navier boundary condition is applied along a dynamic apparent contact angle that accounts for the local interface bending. The robustness of the model is assessed by comparing the toy model results with standard phase field ones on two cases: the steady state profiles of a liquid bridge between two plates moving in opposite directions and the transient behaviors of a drop spreading on a solid with a prescribed equilibrium angle. This offers a practical and efficient alternative to solve contact line problems at lower cost in a sharp interface framework with input parameters from phase field models.
I Introduction
@@
The contact line that forms at the intersection of a fluid-fluid and a solid boundary is a key element of many natural and several technological processes. The dynamics of the contact line has been a subject of debate for several decades [1, 2, 3, 4, 5, 6, 7]. The classical no-slip boundary condition at a moving fluid interface leads to a force singularity at the contact line, as demonstrated Huh and Scriven [8], implying that contact line motion would be impossible.
@@
In the context of continuum mechanics, in sharp-interface models, one common approach is to introduce a small slip velocity at the contact line using the Navier boundary condition. The slip length, typically on the order of nanometers, is often associated with surface roughness or the mean free path of molecules [9]. The Navier slip model combined with an equilibrium contact angle on a mesh-aligned boundary has been extensively studied across various physical setups using different numerical methods. For instance, a rapidly advancing contact line in a curtain coating scenario was simulated in a volume-of-fluid (VOF) framework [10], different numerical methods with the same contact line model were tested in a capillary rise setup [11], a withdrawing plate was simulated using the VOF method with a mesh-dependent slip length [12]. Spreading droplets were explored with mesh-dependent slip [13], and the Level-Set method was employed for a spreading drop case, incorporating contact line hysteresis by varying the contact angle based on the contact line speed [14]. Recently, a Generalized Navier Boundary Condition implemented within the same sharp interface framework was shown to regularize the logarithmic singularity at the contact line by allowing a deviation of dynamic contact angle with respect to the equilibrium one through the uncompensated Young’s stress [15, 16]. These models have in common that they require empirical information on the slip length, which can be viewed as a fitting parameter.
@@
On the other hand, in the phase-field method (PF) the Cahn-Hilliard equations are formulated from the thermodynamics of an immiscible two-component mixture and a phase function is used to represent the moving interface. The velocity field can satisfy a no-slip condition at the contact line, and the contact line moves due the diffusive mass transport that is present in a two-component system [17, 18]. While the phase-field method has the advantage of needing less empirical fitting, it is considerably more expensive. Although it has been shown that it is possible to match the results obtained on dynamic contact lines by fine-tuning the different parameters of the models [19, 20], an ideal model would mimic the mass diffusion by a simple model; that is to obtain results essentially equivalent to the PF ones by doing much cheaper VOF simulations.
@@
The paper is organized as follows. Section II describes the sharp interface framework and Section III the phase field formulation. In Section IV we derive the toy model for the dynamic contact line starting from the phase field approximation of the incompressible two-phase Navier-Stokes equations. In Section V we show a comparison of the toy model VOF results with standard PF ones on two different systems. Conclusion follows in Section VI.
II Sharp interface model
The volume-of-fluid method for representing fluid interfaces coupled with a flow solver is well-known to be suited for solving interfacial flows [21, 22, 23]. In the present work, we use the free software Basilisk, a platform for the solution of partial differential equations on adaptive Cartesian meshes [24, 25, 26]. We consider two incompressible, Newtonian and non-miscible fluids. We assume that there is no mass and heat transfer at the fluid-fluid interface, that the surface tension between the two fluids $\sigma$ is constant, and neglect gravity. The two fluids will be denoted $l$ and $g$ for liquid and gas phases, with densities and dynamical viscosities $\rho_{l}$ and $\rho_{g}$, $\mu_{l}$ and $\mu_{g}$ respectively. In the one-fluid formulation, the unsteady, incompressible, variable density, Navier-Stokes equations read
$\dfrac{\partial\rho}{\partial t}+\mathbf{\nabla}\cdot(\rho\,\mathbf{u})$ $=0,$ (1)
$\rho\left(\dfrac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\mathbf{\nabla}\mathbf{u}\right)$ $=-\mathbf{\nabla}P+\mathbf{\nabla}\cdot(2\mu\mathbf{D})+\mathbf{F}_{\sigma},$ (2)
$\mathbf{\nabla}\cdot\mathbf{u}$ $=0,$ (3)
with $\mathbf{u}\equiv\mathbf{u}(\mathbf{x},t)$ the flow velocity, $P\equiv P(\mathbf{x},t)$ the pressure, $\mathbf{D}=(\mathbf{\nabla}\mathbf{u}+\mathbf{\nabla}\mathbf{u}^{T})/2$ the rate of deformation tensor, $\rho\equiv\rho(\mathbf{x},t)$ the density, $\mu\equiv\mu(\mathbf{x},t)$ the viscosity and $\mathbf{F}_{\sigma}$ the surface tension force. The surface tension term is modeled using the continuum surface force [27] so that $\mathbf{F}_{\sigma}=\sigma\,\kappa\,\nabla\chi$ with $\kappa$ the curvature, $\mathbf{n}$ the normal unit vector to the interface and $\chi$ the Heaviside function
$\chi(\mathbf{x},t)=\left\{\begin{array}[]{rl}0,&\text{in the gas phase},\\ 1,&\text{in the liquid phase},\end{array}\right.$ (4)
also used to locate the interface and define the viscosity and density [23]. The volume fraction $c(\mathbf{x},t)$ corresponding to the average value of $\chi$ in each cell is used to replace the advection of the density (1) by the following advection equation:
$\dfrac{\partial c}{\partial t}+\mathbf{u}\cdot\mathbf{\nabla}c=0.$ (5)
@@ II Sharp interface model
The incompressible Navier-Stokes equations are approximated using a time-staggered approximate projection method on a Cartesian grid. The advection term is discretized with the explicit and conservative Bell–Colella–Glaz second-order unsplit upwind scheme [28]. For the discretization of the viscous diffusion term, a second-order Crank–Nicholson fully-implicit scheme is used. Spatial discretization is achieved using a quadtree (octree in three dimensions) adaptive mesh refinement on collocated grids. The adaptive mesh technique is based on the wavelet decomposition of the variables such as the velocity field or the volume fraction and is efficient and critical to the success of the simulations as it allows high resolution only where needed, therefore reducing the cost of calculation.
@@
The resolution of the surface tension term is directly dependent on the accuracy of the curvature calculation. The Height-Function methodology is a VOF-based technique for calculating interface normals and curvatures [29, 30]. About each interface cell, fluid ‘heights’ are calculated by summing fluid volume in the grid direction closest to the normal of the interface. In two dimensions, a $7\times 3$ stencil around an interface cell is constructed and the heights are evaluated by summing volume fractions horizontally
$h_{j}=\sum_{k=i-3}^{k=i+3}c_{j,k}\Delta_{\text{V}},$ (7)
@@
$\begin{array}[]{c}{\mathbf{n}=\left(h_{x},-1\right)},\\ \\ {\kappa=\dfrac{h_{xx}}{\left(1+h_{x}^{2}\right)^{3/2}}},\end{array}$ (8)
where $h_{x}$ and $h_{xx}$ are discretized using second-order central differences. This method is also used to model the contact angle at the triple point. The orientation of the interface, characterized by the contact angle – the angle between the normal to the interface at the contact line and the normal to the solid boundary – is imposed in the contact line cell. It is important to note that a numerical specification of the contact angle affects the overall flow calculation in two ways: (i) it defines the orientation of the VOF reconstruction in cells that contain the contact line; (ii) it influences the calculation of the surface tension term by affecting the curvature computed in cells at and near the contact line. The boundary conditions in the VOF toy model, derived in Section IV, will result in a simple slip model coupled to a dynamic contact angle. The slip model, or Navier boundary condition, has been implemented and used in various studies within this numerical framework [10, 19]. In two dimensions, the Navier boundary condition reads
$\begin{split}u_{x}|_{y=0}+\lambda\dfrac{\partial u_{x}}{\partial y}|_{y=0}&=U,\\ u_{y}|_{y=0}&=0,\end{split}$ (9)
with $u_{x}|_{y=0}$ and $u_{y}|_{y=0}$ the $x$ and $y$ component of the velocity at the solid boundary, $\lambda$ the slip length, and $U$ the prescribed velocity of the moving solid. Note that (9) is written with the boundary located at $y=0$; the location of the computational boundary will be lifted in the toy model (Section IV) such that the boundary is located at $y=+\lambda$. Moreover, it is important to remark that the accuracy of the numerical method is dependent on the resolution of the smallest length scale, the slip length $\lambda$. In the following numerical setups, the grid size will be chosen such that $\lambda/\Delta_{\text{V}}\simeq 4$.
III Diffuse interface model
In the phase-field method, the governing equations are derived from the thermodynamic potentials of the system, together with the assumption of a surface energy associated with an interface [31, 32, 18]. It is, therefore, possible to consider different physical situations with relative ease. It is also straightforward to implement numerically, since interfaces are not tracked explicitly. Instead, a variable is introduced that has different constant values in the two phases with a steep transition between the two in the diffuse interface [32, 33]. One of the major drawbacks of this method, however, is that the width of the interface $\epsilon$ must be small to match the proper interface dynamics.
@@ III Diffuse interface model
$\dfrac{\partial C}{\partial t}=\nabla\cdot(-\mathbf{F}_{d}-\mathbf{F}_{c}),$ (10)
@@
$\mathbf{F}_{c}=\mathbf{u}\cdot\nabla C,$ (11)
@@
$\phi=\beta\Psi^{\prime}\left(C\right)-\alpha\nabla^{2}C.$ (13)
@@
$\sigma=\dfrac{2}{3}\sqrt{2\alpha\beta},\quad\quad\epsilon=\sqrt{\dfrac{\alpha}{\beta}}.$ (14)
@@
$\Psi\left(C\right)=\dfrac{\left(C+1\right)^{2}\left(C-1\right)^{2}}{4}.$ (15)
With the definitions above (10)- (15), the convection-diffusion equation, referred to as the Cahn-Hilliard (1958b) equation, can be written as
$\dfrac{\partial C}{\partial t}=\nabla\cdot\left[M\nabla\left(\beta\Psi^{\prime}\left(C\right)-\alpha\nabla^{2}C\right)\right]-\mathbf{u}\cdot\nabla C.$ (16)
Equation (16) is the phase-field analog of (5). The boundary conditions for this fourth-order partial differential equation will be detailed in our study of dynamic wetting by the phase-field method. Taking the same notation as in (2), the momentum equation in the Navier-Stokes equations becomes
$\rho\left(\partial_{t}\mathbf{u}+\mathbf{u}\cdot\nabla\mathbf{u}\right)=-\nabla p+\nabla\cdot(2\mu\mathbf{D})+\phi\nabla C\,$ (17)
@@
$\rho({C})$ $\equiv\rho_{l}\dfrac{C+1}{2}+\rho_{g}\dfrac{C-1}{2},$ (18)
@@
We now state the boundary conditions in the presence of a contact line on a solid, characterized by its equilibrium angle (the angle between the fluid-fluid interface and the wall). The convection-diffusion equation (16) is a fourth-order partial differential equation and requires two boundary conditions. First, we impose a non-equilibrium wetting condition [32, 35] on the solid wall,
$-\mu_{f}\epsilon\left(\dfrac{\partial C}{\partial t}+\mathbf{u}\cdot\nabla C\right)=\alpha\nabla C\cdot\mathbf{n}-\sigma\cos\theta_{\mathrm{e}}\>g^{\prime}\left(C\right),$ (19)
@@
$g\left(C\right)=\dfrac{1}{4}(2-3C+C^{3}),$ (20)
@@
$\nabla\phi\cdot\mathbf{n}=0.$ (21)
The phase-field equations are written into the weak form, and solved using an open-source finite-element software FreeFEM by Hecht, Pironneau and others [36, 37], allowing the specification of a finite-element weak form as input. Linear elements were used for the PF variables, while the fluid flow was resolved using Taylor–Hood elements (quadratic for velocity and linear for pressure). The numerical code is available in the Github repository FreeFEM-NS-CH. Similarly to the VOF implementation, the accuracy of the method depends on the resolution of the smallest length scale of the system, the interface width $\epsilon$. In the PF setups, the grid size will be chosen such that $\epsilon/\Delta_{\text{P}}\simeq 16$.
IV The toy model: relating mass flux to curvature